# Utility-Learning Tension in Self-Modifying Agents

### v1

@@ 4 Representational Self-Modification ($\mathcal{M}_{H}$)
$H_{t+1}=\Phi_{H}(H_{t},D_{t},\theta_{t}).$
Data follow $(x,y)\sim\mathcal{D}$ i.i.d.; the training set $S\sim\mathcal{D}^{m}$ and validation set $V\sim\mathcal{D}^{n_{v}}$ are independent. Loss $\ell\in[0,1]$; risk $R(h)=\mathbb{E}[\ell(h(x),y)]$; empirical risks $\widehat{R}_{S}$, $\widehat{R}_{V}$. Capacity is measured by $\mathrm{VC}(\cdot)$. A utility $u$ is reasonable if it is computable, normalized to $[0,1]$, and (U1) non-decreasing in empirical fit on the active finite evidence, and (U2) strictly increasing in a computable capacity bonus $g(\mathrm{VC})$ with $g^{\prime}>0$. The decision rule executes an edit only when an immediate utility increase is formally provable. Proofs are deferred to Appendix 12.
@@
[…]
@@
$\mathcal{H}_{\mathrm{reach}}(u)=\{\,$ $H^{\prime}:\ \exists t\ \text{along a policy-reachable}$
$\text{trajectory under }u\ \text{with }H_{t}=H^{\prime}\,\}.$
Unbounded representational power (URP).
$(\mathcal{H},\Phi_{H})$ has URP if for every $m\in\mathbb{N}$ there exist $H,\theta,D$ with
$\mathrm{VC}\!\big(\Phi_{H}(H,D,\theta)\big)\ \geq\ m.$
Local-URP: for every $H$ there exists an edit with $\mathrm{VC}(\Phi_{H}(H,\cdot,\cdot))\geq\mathrm{VC}(H)+1$ that can fit the current finite sample.
Policy-level learnability boundary
Under (A1–A7), distribution-free learnability is preserved under representational self-modification iff
$\sup_{H^{\prime}\in\mathcal{H}_{\mathrm{reach}}(u)}\mathrm{VC}(H^{\prime})<\infty.$
Sketch. ($\Leftarrow$) A uniform capacity cap gives uniform convergence on a fixed capped reference family for all steps; ERM/AERM yields the standard VC rate for the terminal predictor. ($\Rightarrow$) If capacities along a reachable subsequence diverge, VC lower bounds preclude any distribution-free sample complexity. (Full proof in Appendix 12.)
Two-Gate finite-sample safety
Given $S$ ($|S|=m$) and independent $V$ ($|V|=n_{v}$), a candidate edit producing $H_{\mathrm{new}}$ is accepted only if
(Validation) $\quad\widehat{R}_{V}(h_{\mathrm{new}})\leq\widehat{R}_{V}(h_{\mathrm{old}})-(2\varepsilon_{V}+\tau),$
(Capacity) $\quad h_{\mathrm{new}}\in\mathcal{G}_{K(m)}\ \ \text{with}\ \ \mathrm{VC}(\mathcal{G}_{K(m)})\leq K(m),$
where $\varepsilon_{V}$ is chosen so that, with probability $\geq 1-\delta_{V}$ over $V$,
$\sup_{h\in\mathcal{G}_{K(m)}}\big|R(h)-\widehat{R}_{V}(h)\big|\leq\ \varepsilon_{V}$
$\ (\varepsilon_{V}\asymp\sqrt{(K(m)+\log(1/\delta_{V}))/n_{v}}).$
Then with probability at least $1-\delta_{V}-\delta$ over draws of $V$ and $S$: (i) each accepted edit decreases true risk by at least $\tau$; and
$R(h_{T})\ \leq\ \inf_{h\in\mathcal{G}_{K(m)}}R(h)\ +\ \tilde{O}\!\Big(\sqrt{(K(m)+\log(1/\delta))/m}\Big).$
Validation reuse (fixed ex ante).
The same validation set $V$ may be reused adaptively across many edits provided the capped reference family $\mathcal{G}_{K(m)}$ is fixed before seeing $V$ and the gate thresholds $K(m),\varepsilon_{V},\tau$ do not depend on $V$. If any of these are tuned using $V$, a fresh split or reusable holdout is required (see Appendix 12).
Figure 1: Representational axis $\mathcal{M}_{H}$. TwoGate accepts a few early edits and then plateaus at a lower test loss, while destructive policies continue modifying the hypothesis class and exhibit worsening generalization as complexity increases.
@@
All oracle inequalities are stated on the intersection of two events: (E1) the uniform validation event on $\mathcal{G}_{K(m)}$ (probability $\geq 1-\delta_{V}$) and (E2) the training-side uniform convergence event (probability $\geq 1-\delta$). By a union bound, the final probability is $\geq 1-\delta_{V}-\delta$ and does not depend on the number of accepted edits, since the bound is uniform over the fixed capped family.
@@
Under URP, utilities that reward empirical fit and even a slight increase in capacity can drive $\mathrm{VC}$ unbounded and destroy distribution-free learnability; see Appendix 12.
5 Architectural Self-Modification ($\mathcal{M}_{Z}$)
Setting and reduction (fixed vs. modifiable).
We analyze architectural edits while holding the learning algorithm $A$, substrate $F$, and metacognitive rule $M$ fixed. An architecture $Z\in\mathcal{Z}$ induces a hypothesis class $H(Z)\subseteq\mathcal{Y}^{\mathcal{X}}$. At time $t$, an architectural edit produces
$Z_{t+1}\;=\;\Phi_{Z}(Z_{t},D_{t},\vartheta_{t}),$
$h_{S}(Z_{t+1})\in\underset{h\in H(Z_{t+1})}{\operatorname{argmin}}\widehat{R}_{S}(h)\ \$
$\hskip 14.22636pt\text{(ERM in the induced class)}.$
@@ 5 Architectural Self-Modification ($\mathcal{M}_{Z}$)
$\mathcal{Z}_{\mathrm{reach}}(u)$ $=\big\{\,Z^{\prime}:\exists t\text{ on some proof-triggered}$
$\qquad\text{trajectory from }Z_{0}\text{ under }u\text{ with }Z_{t}=Z^{\prime}\,\big\},$
$\mathcal{H}^{Z}_{\mathrm{reach}}(u)$ $=\{\,H(Z):Z\in\mathcal{Z}_{\mathrm{reach}}(u)\,\}.$
@@
The boundary and Two-Gate guarantees depend only on the capacity of the reachable family, not on explicit capacity rewards: even if $u$ has no bonus term, any policy that permits capacity-increasing edits can cross the boundary unless a cap (e.g., via $K(m)$) is enforced. The stronger “bonus” variant is used only for the destruction theorems.
Every run of $\mathcal{M}_{Z}$ therefore corresponds to a run of $\mathcal{M}_{H}$ over the induced family $\mathcal{H}^{Z}_{\mathrm{reach}}(u)$.
Architectural $\to$ representational reduction
For any reasonable $u$, every proof-triggered trajectory $Z_{0}\!\to\!Z_{1}\!\to\!\cdots$ induces a representational trajectory $H(Z_{0})\!\to\!H(Z_{1})\!\to\!\cdots$ over the fixed reference family $\mathcal{G}$, with ERM/AERM inside each accepted $H(Z_{t})$. Consequently $\mathcal{H}^{Z}_{\mathrm{reach}}(u)=\{H(Z):Z\in\mathcal{Z}_{\mathrm{reach}}(u)\}\subseteq\mathcal{G}$. Proof sketch. The decision semantics and utility are unchanged by renaming states from $Z$ to the induced $H(Z)$.
Architectural boundary via induced policy-reachable family
For any reasonable $u$, distribution-free PAC learnability under architectural self-modification is preserved iff
$\sup_{Z\in\mathcal{Z}_{\mathrm{reach}}(u)}\mathrm{VC}\!\big(H(Z)\big)\ \leq\ K\;<\ \infty.$
Equivalently, preservation holds iff
$\sup_{H^{\prime}\in\mathcal{H}^{Z}_{\mathrm{reach}}(u)}\mathrm{VC}(H^{\prime})\leq K.$
@@
Immediate by reduction to §4: apply the sharp boundary theorem for representation (Thm. 4) to the induced set $\mathcal{H}^{Z}_{\mathrm{reach}}(u)$. ∎
@@
Fix a single parameterized super-family $\mathcal{G}\subseteq\mathcal{Y}^{\mathcal{X}}$ that contains every induced class: $H(Z)\subseteq\mathcal{G}$ for all $Z$. For $K\in\mathbb{N}$ define the proxy-cap subfamily
$\mathcal{G}^{\mathrm{proxy}}_{K}\ :=\ \{\,h\in\mathcal{G}:\ \exists Z\text{ with }B(Z)\leq K$
$\text{ and }h\in H(Z)$ $\,\},$
@@
Two-gate safety for architecture
[…]
Validation reuse. We fix $\mathcal{G}^{\mathrm{proxy}}_{K(m)}$, $K(m)$, and thresholds ex ante (before seeing $V$). If any choice is tuned on $V$, we use a fresh split or a reusable-holdout mechanism; all theorems apply to the fixed family (Appendix 13.3).
Local architectural edits (Local-URPZ).
We say $\mathcal{M}_{Z}$ has Local-URPZ if for every $Z$ there exists a computable edit $(\vartheta,D)$ such that $\mathrm{VC}\!\big(H(\Phi_{Z}(Z,D,\vartheta))\big)\geq\mathrm{VC}(H(Z))+1$ and the new class $H(\Phi_{Z}(Z,D,\vartheta))$ can interpolate the current finite evidence (e.g., fit $S$).
Robust destruction under Local-URPZ
Assume Local-URPZ. For any reasonable utility $u$ (non-decreasing in empirical fit on the active evidence and strictly increasing in a computable capacity bonus $g(\mathrm{VC})$), there exist a distribution $\mathcal{D}$ and sample size $m$ such that the proof-trigger repeatedly accepts local architectural edits that increase capacity, the induced reachable set $\mathcal{H}^{Z}_{\mathrm{reach}}(u)$ has unbounded VC, and distribution-free PAC learnability fails. Sketch. Each local step strictly increases $u$ (fit non-worse + $g$ increases), so the proof-trigger fires. Iteration yields unbounded VC; apply the necessity direction of Thm. 5.
Proxy-cap sufficiency (computable architectural bounds).
Often one has a computable upper bound $B(Z)$ on $\mathrm{VC}(H(Z))$ (e.g., for ReLU networks $B(Z)\!=\!c\,W(Z)\log W(Z)$ with $W(Z)$ parameters).
Proxy-cap two-gate guarantee
If $B(Z)\geq\mathrm{VC}(H(Z))$ for all $Z$ and the capacity gate enforces $B(Z_{\mathrm{new}})\leq K(m)$, then Corollary 5 holds with $\mathcal{G}_{K(m)}=\{\,h\in\mathcal{G}:\ \text{$h$ realizable by some $Z$ with }B(Z)\leq K(m)\,\}$ and with the same rate (the VC bound is applied to $\mathcal{G}_{K(m)}$). Proof. Since $H(Z_{\mathrm{new}})\subseteq\mathcal{G}_{K(m)}$ and $\mathrm{VC}(\mathcal{G}_{K(m)})\leq K(m)$ by construction, the two-gate proof (Appendix 12) applies verbatim. ∎
6 Metacognitive Self-Modification ($\mathcal{M}_{M}$)
Setting (fixed vs. modifiable).
We analyze metacognitive scheduling/filters $M$ while holding representation $H$, architecture $Z$, algorithm $A$, and substrate $F$ fixed; $M$ selects which edits to evaluate and applies acceptance/rejection using finite evidence.
Metacognitive scheduler/filter.
A metacognitive rule $M$ can (i) choose which candidate edit to evaluate, (ii) when/how often to evaluate (scheduling), and (iii) randomize; it then accepts/rejects using only finite evidence (e.g., $S,V$, capacity proxy, edit cost). Let
$\mathcal{H}^{M,H}_{\mathrm{reach}}(u)\$$:=\ \Big\{\,H^{\prime}:\Pr\big[\exists t\ \text{on a proof-triggered}$
$\qquad\text{trajectory from }H_{0}\text{ under }u\text{ filtered}$
$\qquad\text{by }M\text{ with }H_{t}=H^{\prime}\big]>0\,\Big\}.$
@@ 6 Metacognitive Self-Modification ($\mathcal{M}_{M}$)
This is the M-filtered policy-reachable family. (Randomized $M$ is allowed; all guarantees below hold almost surely over $M$’s internal randomness.)
Boundary under metacognitive modifications.
@@
For any reasonable $u$ and metacognitive filter/scheduler $M$, distribution-free PAC learnability is preserved iff
$\sup_{H^{\prime}\in\mathcal{H}^{M,H}_{\mathrm{reach}}(u)}\mathrm{VC}(H^{\prime})\ \leq\ K\;<\ \infty.$
@@
Proof sketch. Apply the representational sharp boundary to the $M$-filtered family $\mathcal{H}^{M,H}_{\mathrm{reach}}(u)$; necessity follows by the same destruction argument when the supremum is unbounded (Appendix 12).
Two-Gate as metacognition (safety + rate).
@@
If $M$ accepts only when (i) $\widehat{R}_{V}(h_{t+1})\leq\widehat{R}_{V}(h_{t})-(2\varepsilon_{V}+\tau)$ and (ii) $\mathrm{VC}(H_{t+1})\leq K(m)$, then with probability $\geq 1-\delta_{V}-\delta$ each accepted edit reduces true risk by $\geq\tau$, and
$R(h_{T})\ \leq\ \inf_{h\in\mathcal{G}_{K(m)}}R(h)\ +\ \tilde{O}\!\Big(\sqrt{\tfrac{K(m)+\log(1/\delta)}{m}}\Big),$
by Thm. 4.
Restorative metacognition (turning destructive utilities safe).
@@
Suppose the unfiltered reachable family satisfies $\sup_{H^{\prime}\in\mathcal{H}_{\mathrm{reach}}(u)}\mathrm{VC}(H^{\prime})=\infty$ (destructive). There exists a metacognitive rule $M$ (e.g., Two-Gate with any computable nondecreasing schedule $K(m)$) such that $\sup_{H^{\prime}\in\mathcal{H}^{M,H}_{\mathrm{reach}}(u)}\mathrm{VC}(H^{\prime})\leq K<\infty$, hence learnability is preserved.
Proof sketch. $M$ simply rejects any proposal that violates the validation margin or the cap; the filtered family is contained in $\mathcal{G}_{K(m)}$ (Appendix 12) and the boundary theorem applies.
Edit efficiency under metacognitive margins.
@@
Under the Two-Gate rule with margin $\tau>0$, along any accepted trajectory
$\#\{\text{accepted edits}\}$ $\leq\frac{R(h_{0})-R^{*}}{\tau},$
$\text{where}\quad R^{*}$ $=\inf_{h\in\mathcal{G}_{K(m)}}R(h).$
@@
Proof. Each accepted edit decreases true risk by at least $\tau$; telescope from $R(h_{0})$ to the floor $R^{*}$.
@@
$M$ may be deterministic or randomized and may schedule when candidate edits are proposed. Because our deviation bound is uniform over the capped family, the safety and rate guarantees hold for every realized trajectory filtered by $M$; scheduling affects efficiency, not the boundary itself.
7 Algorithmic Self-Modification ($\mathcal{M}_{A}$)
Setting (fixed vs. modifiable).
We analyze algorithmic self-modification while holding the hypothesis class $H$, architecture $Z$, substrate $F$, and metacognitive rule $M$ fixed. The learner may change its learning algorithm or schedule (optimizer, step sizes, noise level, stopping rule, etc.), producing an updated training procedure $A_{t+1}=\Phi_{A}(A_{t},D_{t},\vartheta_{t})$. Given a training set $S$ of size $m$, the output predictor is $\hat{h}=\mathsf{Alg}(A,S,H)\in H$.

### v2

@@ 4 Representational Self-Modification $\mathcal{M}_{H}$
$H_{t+1}=\Phi_{H}\!\left[H_{t},E_{t},\theta_{t}\right].$
Data follow $\langle x,y\rangle\sim\mathcal{D}$ i.i.d. The training set $S\sim\mathcal{D}^{m}$ and validation set $V\sim\mathcal{D}^{n_{v}}$ are independent. Loss $\ell\in\left[0,1\right]$. Risk is $R\!\left[h\right]$ and empirical risks are $\widehat{R}_{S}$ and $\widehat{R}_{V}$. Capacity is measured by $\mathrm{VC}\!\left[\cdot\right]$. Proofs are deferred to Appendix 12.
Utility assumptions.
We use two tiers of assumptions on utility. A utility $u$ is reasonable if it is computable from finite state and evidence, normalized to $\left[0,1\right]$, and is non-decreasing in empirical fit on the active finite evidence. For the destruction results only, we additionally assume a capacity-bonus reasonable utility: $u$ strictly increases with a computable bonus term $g\!\left[\mathrm{VC}\right]$ with $g^{\prime}>0$.
@@
We work with a fixed capped reference family $\mathcal{G}_{K\!\left[m\right]}\subseteq\mathcal{Y}^{\mathcal{X}}$, satisfying $\mathrm{VC}\!\left[\mathcal{G}_{K\!\left[m\right]}\right]\leq K\!\left[m\right]$ and fixed ex ante before seeing $V$.
@@
$\mathcal{H}_{\mathrm{reach}}\!\left[u\right]=\left\{H^{\prime}:\ \exists t\ \text{along a policy-reachable trajectory under }u\ \text{with }H_{t}=H^{\prime}\right\}.$
Unbounded representational power.
The pair $\langle\mathcal{H},\Phi_{H}\rangle$ has unbounded representational power if for every $m\in\mathbb{N}$ there exist $H,\theta,E$ with
$\mathrm{VC}\!\left[\Phi_{H}\!\left[H,E,\theta\right]\right]\ \geq\ m.$
Local unbounded representational power holds if for every $H$ there exists an edit that increases VC by at least one and can fit the current finite evidence.
Theorem 1. Policy-level learnability boundary
Under Assumptions A1 through A7, distribution-free learnability is preserved under representational self-modification if and only if
$\sup_{H^{\prime}\in\mathcal{H}_{\mathrm{reach}}\!\left[u\right]}\mathrm{VC}\!\left[H^{\prime}\right]<\infty.$
Sketch. Sufficiency follows because a uniform capacity cap gives uniform convergence on a fixed capped reference family for all steps, and ERM or AERM yields the standard VC rate for the terminal predictor. Necessity follows because if capacities along a reachable subsequence diverge, VC lower bounds preclude any distribution-free sample complexity. Full proof appears in Appendix 12.
Theorem 2. Two-Gate finite-sample safety
Given $S$ with $\lvert S\rvert=m$ and independent $V$ with $\lvert V\rvert=n_{v}$, a candidate edit producing $H_{\mathrm{new}}$ is accepted only if
Validation $\widehat{R}_{V}\!\left[h_{\mathrm{new}}\right]\leq\widehat{R}_{V}\!\left[h_{\mathrm{old}}\right]-\left[2\varepsilon_{V}+\tau\right],$
Capacity $h_{\mathrm{new}}\in\mathcal{G}_{K\!\left[m\right]}\ \ \text{and}\ \ \mathrm{VC}\!\left[\mathcal{G}_{K\!\left[m\right]}\right]\leq K\!\left[m\right],$
where $\varepsilon_{V}$ is chosen so that, with probability at least $1-\delta_{V}$ over $V$,
$\sup_{h\in\mathcal{G}_{K\!\left[m\right]}}\left|R\!\left[h\right]-\widehat{R}_{V}\!\left[h\right]\right|\leq\ \varepsilon_{V},\qquad\varepsilon_{V}\asymp\sqrt{\frac{K\!\left[m\right]+\log\!\left[1/\delta_{V}\right]}{n_{v}}}.$
Then with probability at least $1-\delta_{V}-\delta$ over draws of $V$ and $S$: each accepted edit decreases true risk by at least $\tau$, and
$R\!\left[h_{T}\right]\ \leq\ \inf_{h\in\mathcal{G}_{K\!\left[m\right]}}R\!\left[h\right]\ +\ \tilde{O}\!\left[\sqrt{\frac{K\!\left[m\right]+\log\!\left[1/\delta\right]}{m}}\right].$
Validation reuse, fixed ex ante.
The same validation set $V$ may be reused adaptively across many edits provided the capped reference family $\mathcal{G}_{K\!\left[m\right]}$ is fixed before seeing $V$ and the gate thresholds $K\!\left[m\right]$, $\varepsilon_{V}$, and $\tau$ do not depend on $V$. If any of these are tuned using $V$, a fresh split or a reusable holdout is required, as detailed in Appendix 12.
@@
All oracle inequalities are stated on the intersection of two events: the uniform validation event on $\mathcal{G}_{K\!\left[m\right]}$ with probability at least $1-\delta_{V}$ and the training-side uniform convergence event with probability at least $1-\delta$. By a union bound, the final probability is at least $1-\delta_{V}-\delta$ and does not depend on the number of accepted edits, since the bound is uniform over the fixed capped family.
@@
Under unbounded representational power, utilities that reward empirical fit and even a slight increase in capacity can drive VC unbounded and destroy distribution-free learnability. See Appendix 12.
5 Architectural Self-Modification $\mathcal{M}_{Z}$
Setting and reduction, fixed versus modifiable.
We analyze architectural edits while holding the learning algorithm $A$, substrate $F$, and metacognitive rule $M$ fixed. An architecture $Z\in\mathcal{Z}$ induces a hypothesis class $H\!\left[Z\right]\subseteq\mathcal{Y}^{\mathcal{X}}$. At time $t$, an architectural edit produces
$Z_{t+1}=\Phi_{Z}\!\left[Z_{t},E_{t},\vartheta_{t}\right],\qquad h_{S}\!\left[Z_{t+1}\right]\in\arg\min_{h\in H\!\left[Z_{t+1}\right]}\widehat{R}_{S}\!\left[h\right].$
@@ 5 Architectural Self-Modification $\mathcal{M}_{Z}$
$\mathcal{Z}_{\mathrm{reach}}\!\left[u\right]=\left\{Z^{\prime}:\exists t\ \text{on a proof-triggered trajectory from }Z_{0}\ \text{under }u\ \text{with }Z_{t}=Z^{\prime}\right\},$
and
$\mathcal{H}^{Z}_{\mathrm{reach}}\!\left[u\right]=\left\{H\!\left[Z\right]:Z\in\mathcal{Z}_{\mathrm{reach}}\!\left[u\right]\right\}.$
@@
The boundary and Two-Gate guarantees depend only on the capacity of the reachable family, not on explicit capacity rewards. Even if $u$ has no bonus term, any policy that permits capacity-increasing edits can cross the boundary unless a cap such as $K\!\left[m\right]$ is enforced. The stronger capacity-bonus variant is used only for destruction results.
Every run of $\mathcal{M}_{Z}$ corresponds to a run of $\mathcal{M}_{H}$ over the induced family $\mathcal{H}^{Z}_{\mathrm{reach}}\!\left[u\right]$.
Lemma 3. Architectural to representational reduction
For any reasonable $u$, every proof-triggered trajectory $Z_{0}\to Z_{1}\to\cdots$ induces a representational trajectory $H\!\left[Z_{0}\right]\to H\!\left[Z_{1}\right]\to\cdots$ over a fixed reference family, with ERM or AERM inside each accepted $H\!\left[Z_{t}\right]$. Consequently $\mathcal{H}^{Z}_{\mathrm{reach}}\!\left[u\right]=\left\{H\!\left[Z\right]:Z\in\mathcal{Z}_{\mathrm{reach}}\!\left[u\right]\right\}$. Proof sketch. The decision semantics and utility are unchanged by renaming states from $Z$ to the induced $H\!\left[Z\right]$.
Theorem 4. Architectural boundary via induced reachable family
For any reasonable $u$, distribution-free PAC learnability under architectural self-modification is preserved if and only if
$\sup_{Z\in\mathcal{Z}_{\mathrm{reach}}\!\left[u\right]}\mathrm{VC}\!\left[H\!\left[Z\right]\right]\ \leq\ K\ <\ \infty.$
Equivalently, preservation holds if and only if $\sup_{H^{\prime}\in\mathcal{H}^{Z}_{\mathrm{reach}}\!\left[u\right]}\mathrm{VC}\!\left[H^{\prime}\right]\leq K$.
@@
Immediate by reduction to Section 4. Apply the representational boundary theorem, Theorem 4, to the induced set $\mathcal{H}^{Z}_{\mathrm{reach}}\!\left[u\right]$. ∎
@@
Fix a single parameterized super-family $\mathcal{G}\subseteq\mathcal{Y}^{\mathcal{X}}$ that contains every induced class: $H\!\left[Z\right]\subseteq\mathcal{G}$ for all $Z$. For $K\in\mathbb{N}$ define the proxy-cap subfamily
$\mathcal{G}^{\mathrm{proxy}}_{K}=\left\{h\in\mathcal{G}:\ \exists Z\ \text{with }B\!\left[Z\right]\leq K\ \text{and}\ h\in H\!\left[Z\right]\right\},$
where $B\!\left[Z\right]$ is a computable architectural capacity proxy satisfying $\mathrm{VC}\!\left[H\!\left[Z\right]\right]\leq B\!\left[Z\right]$. Since each accepted $H\!\left[Z\right]$ is a subset of the fixed reference subfamily $\mathcal{G}^{\mathrm{proxy}}_{K}$, capacity is bounded by $\mathrm{VC}\!\left[\mathcal{G}^{\mathrm{proxy}}_{K}\right]\leq K$.
Corollary 5. Two-Gate safety for architecture
Let the capacity gate enforce $H\!\left[Z_{\mathrm{new}}\right]\subseteq\mathcal{G}^{\mathrm{proxy}}_{K\!\left[m\right]}$ with $\mathrm{VC}\!\left[\mathcal{G}^{\mathrm{proxy}}_{K\!\left[m\right]}\right]\leq K\!\left[m\right]$, and the validation gate enforce $\widehat{R}_{V}\!\left[h_{S}\!\left[Z_{\mathrm{new}}\right]\right]\leq\widehat{R}_{V}\!\left[h_{S}\!\left[Z_{\mathrm{old}}\right]\right]-\left[2\varepsilon_{V}+\tau\right]$, where $\varepsilon_{V}$ is chosen by a VC bound on $\mathcal{G}^{\mathrm{proxy}}_{K\!\left[m\right]}$. Then the safety and rate conclusions hold as stated by Theorem 4.
Validation reuse.
We fix $\mathcal{G}^{\mathrm{proxy}}_{K\!\left[m\right]}$, the schedule $K\!\left[m\right]$, and thresholds ex ante before seeing $V$. If any choice is tuned on $V$, we use a fresh split or a reusable holdout mechanism; all theorems apply to the fixed family.
Local architectural edits.
We say $\mathcal{M}_{Z}$ has local unbounded representational power if for every $Z$ there exists a computable edit $\langle\vartheta,E\rangle$ such that $\mathrm{VC}\!\left[H\!\left[\Phi_{Z}\!\left[Z,E,\vartheta\right]\right]\right]\geq\mathrm{VC}\!\left[H\!\left[Z\right]\right]+1$ and the new class can interpolate the current finite evidence.
Proposition 6. Robust destruction under local power
Assume local unbounded representational power for $\mathcal{M}_{Z}$. For any capacity-bonus reasonable utility $u$, there exist a distribution $\mathcal{D}$ and sample size $m$ such that the proof-trigger repeatedly accepts local architectural edits that increase capacity, the induced reachable set $\mathcal{H}^{Z}_{\mathrm{reach}}\!\left[u\right]$ has unbounded VC, and distribution-free PAC learnability fails. Sketch. Each local step strictly increases $u$, so the proof-trigger fires. Iteration yields unbounded VC, then apply Theorem 5.
Proposition 7. Proxy-cap sufficiency
If $B\!\left[Z\right]\geq\mathrm{VC}\!\left[H\!\left[Z\right]\right]$ for all $Z$ and the capacity gate enforces $B\!\left[Z_{\mathrm{new}}\right]\leq K\!\left[m\right]$, then the Two-Gate guarantee holds with
$\mathcal{G}_{K\!\left[m\right]}=\left\{h\in\mathcal{G}:\ \exists Z\ \text{with }B\!\left[Z\right]\leq K\!\left[m\right]\ \text{and}\ h\in H\!\left[Z\right]\right\},$
and with the same VC-rate bound.
6 Metacognitive Self-Modification $\mathcal{M}_{M}$
Setting, fixed versus modifiable.
We analyze metacognitive scheduling and filters $M$ while holding representation $H$, architecture $Z$, algorithm $A$, and substrate $F$ fixed. The rule $M$ selects which edits to evaluate and applies acceptance or rejection using finite evidence.
Metacognitive scheduler or filter.
A metacognitive rule $M$ can choose which candidate edit to evaluate, when to evaluate, and can randomize; it then accepts or rejects using only finite evidence, such as $S$, $V$, a capacity proxy, and edit costs. Let
@@ 6 Metacognitive Self-Modification $\mathcal{M}_{M}$
Randomized $M$ is allowed. All guarantees hold almost surely over the internal randomness of $M$.
Theorem 8. Boundary under metacognitive filtering
@@
For any reasonable $u$ and metacognitive filter $M$, distribution-free PAC learnability is preserved if and only if
$\sup_{H^{\prime}\in\mathcal{H}^{M,H}_{\mathrm{reach}}\!\left[u\right]}\mathrm{VC}\!\left[H^{\prime}\right]\ \leq\ K\ <\ \infty.$
@@
Proof sketch. Apply the representational boundary theorem to the $M$ filtered family. Necessity follows by the same destruction argument when the supremum is unbounded. Full details appear in Appendix 12.
Two-Gate as metacognition.
@@
If $M$ accepts only when the validation margin holds and $\mathrm{VC}\!\left[H_{t+1}\right]\leq K\!\left[m\right]$, then with probability at least $1-\delta_{V}-\delta$ each accepted edit reduces true risk by at least $\tau$, and
$R\!\left[h_{T}\right]\ \leq\ \inf_{h\in\mathcal{G}_{K\!\left[m\right]}}R\!\left[h\right]\ +\ \tilde{O}\!\left[\sqrt{\frac{K\!\left[m\right]+\log\!\left[1/\delta\right]}{m}}\right],$
by Theorem 4.
Restorative metacognition.
@@
Suppose the unfiltered reachable family satisfies $\sup_{H^{\prime}\in\mathcal{H}_{\mathrm{reach}}\!\left[u\right]}\mathrm{VC}\!\left[H^{\prime}\right]=\infty$. There exists a metacognitive rule $M$, such as Two-Gate with any computable nondecreasing schedule $K\!\left[m\right]$, such that $\sup_{H^{\prime}\in\mathcal{H}^{M,H}_{\mathrm{reach}}\!\left[u\right]}\mathrm{VC}\!\left[H^{\prime}\right]\leq K<\infty$, hence learnability is preserved.
Edit efficiency under margins.
@@
Under Two-Gate with margin $\tau>0$, along any accepted trajectory the number of accepted edits is at most
$\frac{R\!\left[h_{0}\right]-R^{\star}}{\tau},\qquad R^{\star}=\inf_{h\in\mathcal{G}_{K\!\left[m\right]}}R\!\left[h\right].$
@@
Each accepted edit decreases true risk by at least $\tau$, so the bound follows by telescoping.
@@
Because the deviation bound is uniform over the capped family, the safety and rate guarantees hold for every realized trajectory filtered by $M$. Scheduling affects efficiency, not the boundary.
7 Algorithmic Self-Modification $\mathcal{M}_{A}$
Setting, fixed versus modifiable.
We analyze algorithmic self-modification while holding the hypothesis class $H$, architecture $Z$, substrate $F$, and metacognitive rule $M$ fixed. The agent may change its update rule, optimizer, or schedule across time, producing
$A_{t+1}=\Phi_{A}\!\left[A_{t},E_{t},\vartheta_{t}\right].$
Given a training set $S$ of size $m$, the output predictor is $\hat{h}=\mathsf{Alg}\!\left[A,S,H\right]\in H$.
Relation to standard online learning.
This axis does not propose a new model class beyond online learning. Any procedure that switches optimizers, such as run SGD then switch to Newton once a condition holds, is still an online learning algorithm because it defines a single map from histories to actions or updates. Our focus is instead a structured subset of online learning algorithms: those that choose among a family of candidate update rules using finite evidence $E_{t}$ to myopically increase utility. This induces an algorithm-selection effect: when the family of reachable update rules has high statistical capacity, the agent can overfit the choice of update rule to $E_{t}$, yielding long-run degradation even when each candidate update rule is itself a valid online learning procedure.